a tank for storing water can be emptied in 10 hours when two faucets are open. if one of the faucet alone will empty the tank in 15 hours, how many hours will it take the faucet alone to empty the tank?
Answer by lwsshak3(11628)

a tank for storing water can be emptied in 10 hours when two faucets are open. if one of the faucet alone will empty the tank in 15 hours, how many hours will it take the faucet alone to empty the tank?
**
1/15=work rate of one faucet
x=hours other faucet will take to empty the tank working alone
1/x=work rate of other faucet
1/10=work rate when 2 faucets are open
..
The sum of individual work rates=work rate when working together
1/15+1/x=1/10
1/x=1/10-1/15=3/30-2/30=1/30
x=30
ans:
hours other faucet will take to empty the tank working alone=30 hrs